OpenStudy (anonymous):

Okay so we're studying Vertical Asymptotes... If f(x)=(x+2)/(x^-2x-3) The asymptotes are 3 and -1, right? Then would the domain be "X is all real numbers except for -1,3"? But on the GDC it only shows 3 as an asymptote...?

OpenStudy (anonymous):

-1 is an asymptote

OpenStudy (anonymous):

So the domain would be "X is all real numbers except for -1 and 3"?

OpenStudy (amistre64):

if there are like factors on top and bottom; they cancel out and create a "jump", or a hole. after canceling out like terms; if there is anything left on the bottom, the value that zeros it out will be an asymptote\[\frac{ab}{bm}\]since the bs cancel, they create a jump discontinuity. After canceling the bs, we are left with an "m" in the bottom which would create an asymptote

OpenStudy (amistre64):

the domain would still exclude b and m regardless

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

but there is nothing to cancel in this problem \[ \frac{x+2}{x^2-2 x-3}=\frac{x+2}{(x-3) (x+1)} \]

OpenStudy (anonymous):

so x=3 and x=-1 are both vertical asymptotes.